A particle moving with K.E. K makes a head on elastic collision with an identical particle at rest.Find the maximum elastic potential energy of the system during collision.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
conserve the komentum first,
mv + 0 = 2mv’
v’ = v/2...........(1)
conservation of enrgy,
0.5mv^2 + 0 = 0.5*2m*v’^2 +  P.E..............(2))
k = 0.5mv^2.........(3)
use all three eqns. and find P.E.